Which function represents the data in the following table of values?
a) x- 0,1,2,3
y- 24,12,6,3
Is it (a)y=24(0.5)^1/x ,(b)y=24(2)^x ,(c) y=12(0.5)^x, or (d) y=24(0.5)^x
b) x-0,3,6,9
y-5,10,20,40
Is it (a) y=5(2)^x ,(b)y=5(3)^x ,(c) y=5(2)^x/3,or (d) y=5(3)^x/2

Answer by stanbon(75887)

a) x- 0,1,2,3
y- 24,12,6,3
Is it (a)y=24(0.5)^1/x ,(b)y=24(2)^x ,(c) y=12(0.5)^x, or (d) y=24(0.5)^x
--------
Notice that when x=0 answer "a" is undefined
answer b = 24 but when x=1 you get 48 instead of 12
answer c = 12
anser d = 24 and you 12 when x = 1; you get 6 when x =2 etc.
So, d is the answer
-------------------------
b) x-0,3,6,9
y-5,10,20,40
Is it (a) y=5(2)^x ,(b)y=5(3)^x ,(c) y=5(2)^x/3,or (d) y=5(3)^x/2
answer a, b, c and d give 5 when x = 0
answer a gives 10 when x=1; the others do not
So, a is the answer
=======================
Cheers,
Stan H.